How many grams of sodium hydroxide (NaOH) are dissolved in a total volume of 100 mL of solution with a sodium hydroxide concentration of 3 mol/L?
1
Expert's answer
2020-09-29T06:56:11-0400

Solution-

For Molarity we will apply formula

Molarity = moles / volume of solution in litre

3 = moles *1000/100

Moles = 3/10 = 0.3 moles

Gm. Of NaOH = 0.3*40

= 12.0gm.